# Personalized Lip Reading: Adapting to Your Unique Lip Movements with Vision and Language

### v1

@@ Abstract
Lip reading aims to predict spoken language by analyzing lip movements. Despite advancements in lip reading technologies, performance degrades when models are applied to unseen speakers due to their sensitivity to variations in visual information such as lip appearances. To address this challenge, speaker adaptive lip reading technologies have advanced by focusing on effectively adapting a lip reading model to target speakers in the visual modality. The effectiveness of adapting language information, such as vocabulary choice, of the target speaker has not been explored in the previous works. Moreover, existing datasets for speaker adaptation have limited vocabulary size and pose variations, limiting the validation of previous speaker-adaptive methods in real-world scenarios. To address these issues, we propose a novel speaker-adaptive lip reading method that adapts a pre-trained model to target speakers at both vision and language levels. Specifically, we integrate prompt tuning and the LoRA approach, applying them to a pre-trained lip reading model to effectively adapt the model to target speakers. In addition, to validate its effectiveness in real-world scenarios, we introduce a new dataset, VoxLRS-SA, derived from VoxCeleb2 and LRS3. It contains a vocabulary of approximately 100K words, offers diverse pose variations, and enables the validation of adaptation methods in wild, sentence-level lip reading for the first time. Through various experiments, we demonstrate that the existing speaker-adaptive method also improves performance in the wild at the sentence level. Moreover, with the proposed adaptation method, we show that the proposed method achieves larger improvements when applied to the target speaker, compared to the previous works.
@@ Introduction
Table 1: Comparison of VoxLRS-SA dataset with publicly available speaker adaptive lip reading datasets. This table outlines the differences in dataset configurations across various dimensions including number of speakers, duration in hours, word count, pose and background diversity, and the presence of transcription labels. The table highlights the comprehensive nature and the enhanced diversity of the VoxLRS-SA dataset.
@@
To address these limitations, we propose a novel method to adapt the model to unseen target speakers in vision and language levels. This methodology comprises of two primary components: 1) For vision level adaptation, the proposed method adjusts the pre-trained lip reading model to adapt to the lip appearance, movements, and speaking speed of the target unseen speaker while minimizing the number of training parameters. 2) For language level adaptation, we aim to adapt the pre-trained model to an individual’s unique linguistic patterns such as frequent vocabulary choice at the language-level, when predicting the target speaker’s spoken language.
@@
Specifically, we effectively design the vision-level adaptation by employing both padding prompts (Kim, Kim, and Ro 2022) and LoRA (He et al. 2024) in the visual encoder different from the previous speaker adaptive lip reading methods solely relying on prompt or LoRA. Moreover, we focus on two components within the visual encoder. Initially, the padding prompt with LoRA is applied to the spatial encoding stage, which is crucial for adapting to the lip appearances of the target speaker. Subsequently, it is applied to the temporal encoding stage, enabling the model to better account for variations in lip movement speeds and styles specific to each speaker, thus enhancing the model’s accuracy in capturing dynamic visual cues. For the language-level adaptation, we apply an input prompt tuning to a decoder by using the fact that input-level modification of pre-trained models on adapting the model on different tasks is effective as evidence (Brown et al. 2020; Zhou et al. 2022). With the input prompt tuning, the decoder learns speaker-specific language modeling probability when predicting the spoken language of the target speaker.
@@
To validate our proposed speaker adaptation method in real-world scenarios, we introduce a new dataset named VoxLRS-SA, derived from VoxCeleb2 (Chung, Nagrani, and Zisserman 2018) and LRS3 (Afouras, Chung, and Zisserman 2018) datasets. Given that both sources consist of about 1700 hours of YouTube videos, they naturally capture diverse speaker poses and contain rich vocabulary. Moreover, the VoxLRS-SA dataset combines the strengths of its parent datasets: LRS3, which includes text transcription labels but lacks specific speaker information, and VoxCeleb2, which provides detailed speaker information but does not include text transcriptions. To address these gaps, we utilize a pre-trained Automatic Speech Recognition (ASR) model (Radford et al. 2023) to generate automatic text labels (Ma et al. 2023; Yeo et al. 2024c) for the VoxCeleb2 dataset. Additionally, we employ a face recognition technology (Serengil and Ozpinar 2024) to generate speaker information for the LRS3 dataset.
@@
The key contributions of this paper are as follows: 1) To the best of our knowledge, this is the first speaker adaptive lip reading method leveraging multimodal information at vision and language levels and exploring its effectiveness in wild sentence-level lip reading. Moreover, the proposed method effectively adapts to a target speaker by employing both prompts and LoRA approaches. 2) We introduce the VoxLRS-SA dataset to validate the speaker adaptive lip reading methods in real-world scenarios. It contains diverse poses and about 100K vocabulary size. 3) Through comprehensive experiments validation on the VoxLRS-SA dataset, we show that the previous speaker-adaptive approach also enhances sentence-level performance in real-world settings. Furthermore, our newly proposed adaptation method surpasses previous methods in improving outcomes when tailored to specific target speakers.
@@
In word-level lip reading, the introduction of publicly available datasets such as LRW (Chung and Zisserman 2017) and LRW-1000 (Yang et al. 2019) has been crucial. Based on these databases, (Stafylakis and Tzimiropoulos 2017) designed 3D convolutional layers and 2D ResNet architectures as the front-end, with LSTM networks as the back-end. Additionally, two-stream networks, which utilize both raw video and optical flow, have been proposed by (Xiao et al. 2020; Weng and Kitani 2019). Some works (Kim, Yeo, and Ro 2022; Yeo, Kim, and Ro 2023) have used the audio signals to enhance the visual representations by learning visual-to-audio mapping in the memory network.
@@
In sentence-level lip reading, the release of the large-scale LRS2 (Afouras et al. 2018) and LRS3 (Afouras, Chung, and Zisserman 2018) databases has been important. With the development of these datasets, the end-to-end model trained using Connectionist Temporal Classification (CTC) loss (Graves et al. 2006) was proposed by (Assael et al. 2016). Subsequently, (Son Chung et al. 2017) developed a model utilizing the Seq2Seq architecture (Sutskever, Vinyals, and Le 2014), and further architectural enhancements were achieved by (Afouras et al. 2018) with the implementation of the Transformer model (Vaswani et al. 2017). Recently, state-of-the-art models (Ma, Petridis, and Pantic 2021) have employed a joint CTC/Attention (Watanabe et al. 2017) loss with the Conformer (Gulati et al. 2020) architecture. (Shi et al. 2022; Haliassos et al. 2022) utilized audio-visual data to train the backbone models in a self-supervised manner, and prove its effectiveness. Recently, with the development of self-supervised learning using audio-visual data, (Kim et al. 2024, 2023) have expanded lip reading technologies to multilingual.
@@
Although there has been great development in lip reading, these systems suffer from degraded performance when applied to unseen speakers not included in the training dataset, due to individual variations in lip appearances and movements. Therefore, it is necessary to explore speaker adaptation techniques to narrow the performance gap between seen and unseen speakers in lip reading. However, existing datasets (Kim, Kim, and Ro 2022; Cooke et al. 2006) for speaker-adaptive lip reading do not cover real-world scenarios due to their limited vocabulary size and lack of pose diversity. To address this issue, we propose a novel dataset, VoxLRS-SA containing about 100K vocabulary size and diverse pose videos.
@@
Figure 1: Overview of the proposed lip reading model architecture illustrating dual adaptation strategies. For the visual level adoption, padding prompt and LoRA are applied to the spatial encoding stage such as the visual front-end, crucial for adapting to each speaker’s unique lip features. Moreover, LoRA also is utilized in the visual back-end to adapt the lip movement speed and style of the target speaker. For language level adaptation, input prompt tuning allows the decoder to learn speaker-specific language probabilities, enhancing its ability to predict the spoken language of the target speaker.
@@
Speaker adaptation (Miao, Zhang, and Metze 2015) was primarily developed for Automatic Speech Recognition (ASR). Early work (Liao, McDermott, and Senior 2013) fine-tuned various parts of the network to adapt the model to the target speaker. Some studies (Li and Sim 2010; Seide et al. 2011) suggested adding additional speaker-dependent layers to the pre-trained model. In the (Swietojanski and Renals 2014), adding speaker-dependent vectors to all pre-trained hidden layers was introduced to adjust them for the test speaker. To provide additional inputs that vary by speaker, (Abdel-Hamid and Jiang 2013b, a; Xue et al. 2014) proposed speaker codes. Recently, these works (Klejch et al. 2019; Huang et al. 2020) have explored the utilization of meta-learning and speech synthesis for speaker adaptation, respectively.
In contrast to the efforts for speaker adaptation in ASR, only a few works handled speaker adaptation for lip reading. Motivated by the traditional speaker adaptation methods (Anastasakos, McDonough, and Makhoul 1997; Gopinath 1998) of ASR, (Almajai et al. 2016) introduced utilize these works into lip reading. Recently, (Kim, Kim, and Ro 2022) proposed on adapting the CNN layer of a visual encoder to the speaker’s unique features such as the shape of the lips by using user-dependent padding. Moreover, (Kim, Kim, and Ro 2023; He et al. 2024) developed the speaker adaptive lip reading by adapting the visual encoder to target the speaker in input, spatial, and temporal levels. (Li et al. 2024; Wu et al. 2024) proposed using lip landmarks to mitigate that lip reading systems are sensitive to individual variation information such as lip appearances.
Past research has concentrated primarily on visual cues for speaker adaptation, neglecting the importance of individual linguistic patterns. Previous works such as (Kolár, yang, and Shriberg 2007; Guz et al. 2010) utilized a speaker’s lexical pattern for dialogue act and sentence segmentation tasks. However, leveraging linguistic cues have not yet been integrated into speaker-adaptive lip reading. By incorporating multimodal information, we improve speaker-specific adaptations, providing a more holistic approach to speaker adaptation.
Methods
@@
Recent work (Yeo et al. 2024a) proves that the LLM’s context modeling capability effectively translates visual speech into sentences. Motivated by this, we employ this framework as our baseline lip reading model. As illustrated in Figure.1, the architecture of the lip reading model based on (Yeo et al. 2024a) consists of two components, a visual encoder for modeling spatio-temporal information from lip movements, and an LLM as a decoder for predicting the sentence by taking visual features as input. To adapt the pre-trained lip reading model to target speakers, we propose a two-fold approach that addresses both the vision and language levels.
@@
Given the speaker-specific training dataset for speaker adaptation $\tau=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{T\times H\times W\times C}$ is the lip movements video of the target unseen speaker, $T$ is the frame length of the video, $y_{i}$ is the text transcriptions corresponding to its lip movements, and $N$ is the number of training samples. To adapt the appearances of the target speaker, the padding prompt and LoRA are employed in the CNN layer of the front-end. These learnable parameters enable the front-end to adapt to the target speaker without any modifications to the pre-trained weights. Specifically, the padding prompt $p$ replaces the zero padding of the CNN layer. At the same time, the CNN layer is reconstructed with the low-rank decomposition matrices. With these two additional parameters, the spatial information encoding process can be formulated as follows: $f_{s}=\mathcal{F}_{\theta,p}(x_{i})$, where $f_{s}\in\mathbb{R}^{T\times D}$ spatial feature of lip movements, $D$ is the embedding dimension of the front-end, the ${\theta}$ indicates the weights of the front-end with the trainable parameters constructing the low-rank decomposition matrices, and $p$ indicates the replacement zero-padding with the padding prompts in the front-end.
For a better modeling of the speaker-specific temporal information such as speaking speed, the back-end needs to be adjusted. In order to effectively adjust without pre-trained weights of the back-end, the LoRA is also utilized in the back-end. Specifically, the LoRA is applied to the attention layer to learn the temporal dependencies between any two positions of speaker-specific lip movements. It can be formulated as follows: $f_{v}=\mathcal{B}_{\theta}(f_{s})$, where $f_{v}\in\mathbb{R}^{T\times D}$ is a visual feature containing spatio-temporal information of lip movements, the $\theta$ indicates the weights of the back-end with the trainable parameters constructing the low-rank decomposition matrices. The visual features encoded from the visual encoder are fed into the projector, to align it with the embedding space of the LLM. These aligned visual features are denoted as $f_{v}^{\prime}\in\mathbb{R}^{T\times D_{L}}$, where $D_{L}$ is the embedding dimension of LLM. With the frozen LLM taking aligned visual features as input, the predicted sentence $\hat{y}_{i}$ is obtained. From this prediction, the cross entropy loss $\mathcal{L}=CE(\hat{y_{i}},y_{i})$ is calculated and, the visual encoder is optimized to adapt to the target speaker at the vision level.
@@
Speaker-specific characteristics encompass not only visual cues such as the appearance of lips but also linguistic patterns, including preferred vocabulary choices. Building on this, language level adaptation focuses on tailoring a pre-trained lip-reading model to align with speaker-specific linguistic patterns, enhancing speaker adaptation. In order to learn these speaker-specific linguistic patterns, our approach utilizes the input prompt tuning motivated by (Brown et al. 2020; Zhou et al. 2022).
Specifically, given the visual feature $f_{v}^{\prime}$ aligned with the embedding space of LLM, the input prompts are concatenated with these features along the temporal dimension. The input prompt has the same dimension as that of the dimension of embedding space of LLM and can be denoted as $p_{i}\in\mathbb{R}^{N_{p}\times D_{L}}$, where $N_{p}$ is the length of prompts. Since the self-attention layers in the LLM, model temporal dependencies between input prompts and visual features, these prompts can influence the LLM’s sentence predictions, It can be formulated as follows: $\hat{y}_{i}=\mathcal{D}(p_{i}\oplus f_{v}^{\prime})$, where the $\oplus$ is the concatenate function. Furthermore, the input prompts are trained to predict more accurate speaker-specific language modeling probabilities using target speaker data. This loss can be formulated as $\mathcal{L}=CE(\hat{y_{i}},y_{i})$, and its loss is minimized to optimize the input prompt for language-level adaptation.
@@
In order to validate the effectiveness of the proposed method in real-world scenarios, we develop a VoxLRS-SA dataset, which contains about 100K sizes of vocabulary, and diversity pose of speakers. This dataset is based on the LRS3 and VoxCeleb2 datasets. While the widely used LRS3 dataset for sentence-level lip reading in the wild contains human-annotated transcription labels, there is no speaker ID information in each video. In contrast to the LRS3 dataset, the VoxCeleb2 datasets have speaker ID information and do not include transcription labels. To address this gap, we generate the speaker ID information in the LRS3 dataset by using face verification (Kim, Kim, and Ro 2022) and make the pseudo-labels in the VoxCeleb2 dataset by employing a pre-trained ASR model (Ma et al. 2023; Yeo et al. 2024c).
@@
Recent work (Ma et al. 2023; Yeo et al. 2024c) has demonstrated that a lip reading model trained using automatically generated transcriptions through a pre-trained ASR model performs similarly to one trained with human-annotated labels. Motivated by these works, we utilize the Whisper (Radford et al. 2023) ASR model to generate automatic labels for the videos classified as the English portion of the VoxCeleb2 dataset. Then, we choose 20 speakers who have more than 50 minutes of video content to construct the test and adaptation (or validation) sets. Additionally, we manually clean the text transcriptions of the test set to accurately evaluate the effectiveness of the speaker adaptation method.
Speaker annotation
Since the pseudo-labels are generated by using the pre-trained ASR model, their transcription quality may be worse than the human-annotated labels. To address this issue, we combine VoxCeleb2 with the LRS3 dataset. Note that the LRS3 dataset provides manually annotated transcriptions corresponding to 433 hours of videos.
However, the LRS3 dataset does not include speaker ID information, and it is necessary to identify the existence of overlapping speakers between the two datasets. To identify overlapping speakers between LRS3 and VoxCeleb2, we first generate the speaker ID information in these two datasets using a method similar to the one described in (Kim, Kim, and Ro 2022), employing a face recognition model. The labeling processes of speaker information are as follows: From each video in LRS3 and VoxCeleb2, we divide the video into three equal segments and extract one face image from each segment. Then, we employ the Facenet (Serengil and Ozpinar 2024) face recognition model to encode the facial images into features of speaker information. To match speakers between the two datasets, we compare the feature vectors from LRS3 and VoxCeleb2. This comparison is done using a similarity measure, such as cosine similarity, to identify the overlapping speakers.
@@
VoxLRS-SA The proposed VoxLRS-SA dataset is divided into two parts corresponding to each purpose. Firstly, the Baseline part of the VoxLRS-SA is designed to train the baseline lip reading model. This training dataset is composed of 1.7K hours of video data from about 9K speakers. Secondly, the Adaptation part is proposed to adapt the pre-trained lip reading model to the target speaker and evaluate the effectiveness of the adaptation method. For this purpose, we construct the Adaptation part with 20 speakers, which do not overlap with the 9K speakers of the Baseline part. The train split contains video data varying from 45 minutes per speaker, with over about 5 minutes of data for each of the 20 speakers in the test and val sets. The detailed data information is shown in Table 2.
VoxCeleb2 The VoxCeleb2 (Chung, Nagrani, and Zisserman 2018) is a speaker verification dataset. It contains only speaker ID information for over 6,000 speakers and 2,442 hours of multilingual videos. By referring to the (Shi et al. 2022), we utilize only an English portion of this dataset, which amounts to 1,326 hours of content.
LRS3 The LRS3 (Afouras, Chung, and Zisserman 2018) is a widely used sentence-level lip reading dataset comprising 433 hours of English talking face videos and human-annotated transcriptions. These talking face videos are collected from TED and TEDx talks, they contain diverse poses.

### v2

@@ Abstract
Lip reading aims to predict spoken language by analyzing lip movements. Despite advancements in lip reading technologies, performance degrades when models are applied to unseen speakers due to their sensitivity to variations in visual information such as lip appearances. To address this challenge, speaker adaptive lip reading technologies have advanced by focusing on effectively adapting a lip reading model to target speakers in the visual modality. However, the effectiveness of adapting language information, such as vocabulary choice, of the target speaker has not been explored in previous works. Additionally, existing datasets for speaker adaptation have limited vocabulary sizes and pose variations, which restrict the validation of previous speaker-adaptive methods in real-world scenarios. To address these issues, we propose a novel speaker-adaptive lip reading method that adapts a pre-trained model to target speakers at both vision and language levels. Specifically, we integrate prompt tuning and the LoRA approach, applying them to a pre-trained lip reading model to effectively adapt the model to target speakers. Furthermore, to validate its effectiveness in real-world scenarios, we introduce a new dataset, VoxLRS-SA, derived from VoxCeleb2 and LRS3. It contains a vocabulary of approximately 100K words, offers diverse pose variations, and enables the validation of adaptation methods in the wild, sentence-level lip reading for the first time in English. Through various experiments, we demonstrate that the existing speaker-adaptive method also improves performance in the wild at the sentence level. Moreover, we show that the proposed method achieves larger improvements compared to the previous works.
Codehttps://bit.ly/Personalized-Lip-Reading
Datasetshttps://bit.ly/VoxLRS-SA-Dataset
Extended versionhttps://bit.ly/2409-00986
@@ Introduction
Table 1: Comparison of our dataset with publicly available speaker adaptive lip reading datasets. This table outlines the differences in dataset configurations across various factors such as the number of words, pose, and background diversity.
@@
To address these limitations, we propose a novel method to adapt the model to unseen target speakers in vision and language levels. This methodology comprises two primary components: 1) For vision level adaptation, the proposed method adjusts the pre-trained lip reading model to adapt to the lip appearance, movements, and speaking speed of the target unseen speaker while minimizing the number of training parameters. 2) For language level adaptation, we aim to adapt the pre-trained model to an individual’s unique linguistic patterns such as frequent vocabulary choice at the language level, when predicting the target speaker’s spoken language.
@@
Specifically, we effectively design the vision-level adaptation by employing both padding prompts (Kim, Kim, and Ro 2022) and LoRA (He et al. 2024) in the visual encoder, different from previous speaker adaptive lip reading methods solely relying on prompt or LoRA. Moreover, we focus on two components within the visual encoder. Initially, the padding prompt with LoRA is applied to the spatial encoding stage, which is crucial for adapting to the lip appearances of the target speaker. Subsequently, it is applied to the temporal encoding stage, enabling the model to better account for variations in lip movement speeds and styles specific to each speaker, thus enhancing the model’s accuracy in capturing dynamic visual cues. For language-level adaptation, we apply input prompt tuning to a decoder based on the fact that input-level modification of pre-trained models are effective for adapting them to different tasks (Brown et al. 2020; Zhou et al. 2022). With input prompt tuning, the decoder learns the probability of language modeling specific to the speaker when predicting the spoken language of the target speaker.
@@
To validate our proposed speaker adaptation method in real-world scenarios, we introduce a new dataset named VoxLRS-SA, derived from VoxCeleb2 (Chung, Nagrani, and Zisserman 2018) and LRS3 (Afouras, Chung, and Zisserman 2018) datasets. Together, these sources consist of about 1700 hours of YouTube videos, and as a result, VoxLRS-SA naturally captures diverse speaker poses and contains a rich vocabulary. Moreover, the VoxLRS-SA dataset combines the strengths of its parent datasets: LRS3, which includes text transcription labels but lacks specific speaker information, and VoxCeleb2, which provides detailed speaker information but does not include text transcriptions. To address these gaps, we utilize a pre-trained Automatic Speech Recognition (ASR) model (Radford et al. 2023) to generate automatic text labels (Ma et al. 2023; Yeo et al. 2024c) for the VoxCeleb2 dataset. Additionally, we employ a face recognition technology (Serengil and Ozpinar 2024) to generate speaker information for the LRS3 dataset.
@@
The key contributions of this paper are as follows: 1) To the best of our knowledge, this is the first speaker adaptive lip reading method leveraging multimodal information at vision and language levels and exploring its effectiveness in the wild sentence-level lip reading. Moreover, the proposed method effectively adapts to a target speaker by employing both prompts and LoRA approaches. 2) We introduce the VoxLRS-SA dataset to validate speaker adaptive lip reading methods in the wild, sentence-level lip reading in English. It contains diverse poses and a vocabulary size of approximately 100K. 3) Through comprehensive experimental validation on the VoxLRS-SA dataset, we show that previous speaker-adaptive approaches also improve sentence-level performance in real-world settings. Furthermore, our newly proposed adaptation method surpasses previous methods in improving outcomes when tailored to specific target speakers.
@@
In word-level lip reading, the introduction of publicly available datasets such as LRW (Chung and Zisserman 2017) and LRW-1000 (Yang et al. 2019) has been crucial. Based on these databases, (Stafylakis and Tzimiropoulos 2017) designed 3D convolutional layers and 2D ResNet architectures as the front-end, with LSTM networks as the back-end. Additionally, two-stream networks, which utilize both raw video and optical flow, have been proposed by (Xiao et al. 2020; Weng and Kitani 2019). Some works (Kim, Yeo, and Ro 2022; Yeo, Kim, and Ro 2023) have used audio signals to enhance the visual representations by learning visual-to-audio mapping in the memory network.
@@
In sentence-level lip reading, the release of the large-scale LRS2 (Afouras et al. 2018) and LRS3 (Afouras, Chung, and Zisserman 2018) databases has been important in English. With the development of these datasets, an end-to-end model trained using Connectionist Temporal Classification (CTC) loss (Graves et al. 2006) was proposed by (Assael et al. 2016). Subsequently, (Son Chung et al. 2017) developed a model utilizing the Seq2Seq architecture (Sutskever, Vinyals, and Le 2014), and further architectural enhancements were achieved by (Afouras et al. 2018) with the implementation of the Transformer model (Vaswani et al. 2017). Recently, state-of-the-art models (Ma, Petridis, and Pantic 2021) have employed a joint CTC/Attention (Watanabe et al. 2017) loss with the Conformer (Gulati et al. 2020) architecture. (Shi et al. 2022; Haliassos et al. 2022) utilized audio-visual data to train backbone models in a self-supervised manner and proved its effectiveness. In the context of other languages, LIP-RTVE dataset (Gimeno-Gómez and Martínez-Hinarejos 2023) provides a resource for automatic lip reading in continuous Spanish in the wild, featuring both speaker-independent and speaker-dependent partitions. Furthermore, with the development of self-supervised learning using audio-visual data, (Kim et al. 2024, 2023) have begun to explore lip reading technologies for multiple languages.
@@
Although there has been great development in lip reading, these systems suffer from degraded performance when applied to unseen speakers not included in the training dataset, due to individual variations in lip appearances and movements. Therefore, it is necessary to explore speaker adaptation techniques to narrow the performance gap between seen and unseen speakers in lip reading. However, existing datasets (Kim, Kim, and Ro 2022; Cooke et al. 2006) for speaker-adaptive lip reading do not cover real-world scenarios for English, due to their limited vocabulary size and lack of pose diversity. To address this issue, we propose a novel dataset VoxLRS-SA, which contains about 100K vocabulary size and diverse pose videos.
@@
Figure 1: Overview of the proposed lip reading model architecture illustrating dual adaptation strategies. Visual level adaptation employs padding prompts and LoRA at the front-end for unique lip features and at the back-end for lip movement speed and style. Language level adaptation uses input prompt tuning to learn the language modeling probability of each speaker.
@@
Speaker adaptation (Miao, Zhang, and Metze 2015) was primarily developed for Automatic Speech Recognition (ASR). An early work (Liao, McDermott, and Senior 2013) fine-tuned various parts of the network to adapt the model to a target speaker. Other studies (Li and Sim 2010; Seide et al. 2011) suggested adding speaker-dependent layers to a pre-trained model. In (Swietojanski and Renals 2014), speaker-dependent vectors were added to all pre-trained hidden layers to adjust them for the test speaker. To provide additional inputs that vary by speaker, (Abdel-Hamid and Jiang 2013b, a; Xue et al. 2014) proposed speaker codes. Recently, works (Klejch et al. 2019; Huang et al. 2020) have explored the utilization of meta-learning and speech synthesis for speaker adaptation, respectively.
In contrast to the efforts for speaker adaptation in ASR, only a few works handled speaker adaptation for lip reading. Motivated by the traditional speaker adaptation methods (Anastasakos, McDonough, and Makhoul 1997; Gopinath 1998) of ASR, (Almajai et al. 2016) introduced utilize these works into lip reading. Recently, (Kim, Kim, and Ro 2022) proposed adapting the CNN layer of a visual encoder to the speaker’s unique features, such as the shape of the lips, by using user-dependent padding. Moreover, (Kim, Kim, and Ro 2023; He et al. 2024) developed speaker adaptive lip reading by adapting the visual encoder to target the speaker at the input, spatial, and temporal levels. (Li et al. 2024; Wu et al. 2024) proposed using lip landmarks to mitigate the sensitivity of lip reading systems to individual variation information such as lip appearances.
Prior work has concentrated primarily on visual cues for speaker adaptation, overlooking the importance of individual linguistic patterns. (Kolár, yang, and Shriberg 2007; Guz et al. 2010) have utilized a speaker’s lexical pattern for dialogue act and sentence segmentation tasks. However, leveraging linguistic cues has not yet been integrated into speaker-adaptive lip reading. By incorporating multimodal information, we improve speaker-specific adaptations, providing a more holistic approach to speaker adaptation.
Method
@@
Recent work (Yeo et al. 2024a) proves that the LLM’s context modeling capability is particularly effective in lip reading systems. Motivated by this, we employ this framework as our baseline lip reading model. As illustrated in Figure.1, the architecture of the lip reading model consists of two components, a visual encoder for modeling spatio-temporal information from lip movements, and an LLM as a decoder for predicting the sentence by taking visual features as input. To adapt the pre-trained lip reading model to target speakers, we propose a two-fold approach that addresses both the vision and language levels.
@@
Given the speaker-specific training dataset for speaker adaptation $\tau=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{T\times H\times W\times C}$ is the lip movements video of the target unseen speaker, $T$ is the frame length of the video, $y_{i}$ is the text transcriptions corresponding to its lip movements, and $N$ is the number of training samples. To adapt the appearances of the target speaker, a padding prompt (Kim, Kim, and Ro 2023) and LoRA (Hu et al. 2021) are employed in the CNN layer of the front-end. These learnable parameters enable the front-end to adapt to the target speaker without any modifications to the pre-trained weights. Specifically, the padding prompt $p$ replaces the zero padding of the CNN layer. At the same time, the CNN layer is reconstructed with the low-rank decomposition matrices. With these two additional parameters, the spatial information encoding process can be formulated as follows: $f_{s}=\mathcal{F}_{\theta,p}(x_{i})$, where $f_{s}\in\mathbb{R}^{T\times D}$ spatial feature of lip movements, $D$ is the embedding dimension of the front-end, ${\theta}$ indicates the weights of the front-end with the trainable parameters constructing the low-rank decomposition matrices, and $p$ indicates the replacement zero-padding with the padding prompts in the front-end.
For a better modeling of the speaker-specific temporal information such as speaking speed, the back-end needs to be adjusted. In order to effectively adjust without pre-trained weights of the back-end, LoRA is also utilized in the back-end. Specifically, LoRA is applied to the attention layer to learn the temporal dependencies between any two positions of speaker-specific lip movements. It can be formulated as follows: $f_{v}=\mathcal{B}_{\theta}(f_{s})$, where $f_{v}\in\mathbb{R}^{T\times D}$ is a visual feature containing spatio-temporal information of lip movements, and $\theta$ indicates the weights of the back-end with the trainable parameters constructing the low-rank decomposition matrices. The visual features encoded from the visual encoder are fed into the projector, to align it with the embedding space of the LLM. These aligned visual features are denoted as $f_{v}^{\prime}\in\mathbb{R}^{T\times D_{L}}$, where $D_{L}$ is the embedding dimension of LLM. With the frozen LLM taking aligned visual features as input, the predicted sentence $\hat{y}_{i}$ is obtained. From this prediction, the cross entropy loss $\mathcal{L}=CE(\hat{y_{i}},y_{i})$ is calculated, and the visual encoder is optimized to adapt to the target speaker at the vision-level.
@@
Speaker-specific characteristics encompass not only visual cues such as the appearance of lips but also linguistic patterns, including preferred vocabulary choices. Since individuals exhibit unique linguistic patterns, the likelihood of specific sequences of words occurring varies from person to person. To learn these speaker-specific linguistic patterns, our approach utilizes an input prompt tuning motivated by (Brown et al. 2020; Zhou et al. 2022; Lester, Al-Rfou, and Constant 2021).
Specifically, given that the visual feature $f_{v}^{\prime}$ is already aligned with the embedding space of the LLM, the input prompts, as trainable parameters, are concatenated with these features along the temporal dimension and then fed into the LLM as input. Consequently, the input prompt is designed to have the same dimension as that of the LLM embedding space dimension and can be denoted as $p_{i}\in\mathbb{R}^{N_{p}\times D_{L}}$, where $N_{p}$ is the length of the prompts. With the self-attention layers in the LLM and model temporal dependencies between input prompts and visual features, these prompts can influence speaker-specific sentence predictions. It can be formulated as follows: $\hat{y}_{i}=\mathcal{D}(p_{i}\oplus f_{v}^{\prime})$, where the $\oplus$ is the concatenate function and $\mathcal{D}$ represents the LLM functioning as the decoder for sentence prediction. In addition, the input prompts are trained to predict more accurate language modeling probabilities for specific speakers using target speaker data. This objective is achieved by minimizing the cross-entropy loss, which can be formulated as $\mathcal{L}=CE(\hat{y_{i}},y_{i})$, which optimizes the input prompt for language-level adaptation.
@@
To validate the effectiveness of the proposed method in real-world scenarios for English, we develop the VoxLRS-SA dataset based on LRS3 and VoxCeleb2 datasets. While the widely used LRS3 dataset for sentence-level lip reading in the wild includes human-annotated transcription labels, it does not contain Speaker ID information for each video. In contrast, the VoxCeleb2 dataset includes Speaker ID information but does not provide text transcriptions. To address this gap, we generate the speaker ID information for the LRS3 dataset by using face verification (Kim, Kim, and Ro 2022) and make pseudo text labels for the VoxCeleb2 dataset by employing a pre-trained ASR model (Ma et al. 2023; Yeo et al. 2024c).
@@
Recent works (Ma et al. 2023; Yeo et al. 2024c) have demonstrated that a lip reading model trained using automatically generated transcriptions through a pre-trained ASR model performs similarly to one trained with human-annotated labels. Motivated by these works, we utilize the Whisper (Radford et al. 2023) ASR model to generate automatic labels for the videos classified as the English portion of the VoxCeleb2 dataset. Then, we choose 20 speakers who have more than 50 minutes of video content to construct the test and adaptation (or validation) sets. Additionally, we manually clean the text transcriptions of the test set to accurately evaluate the effectiveness of the speaker adaptation method.
Speaker ID information annotation
Since the pseudo-labels are generated using the pre-trained ASR model, their transcription quality may be worse than human-annotated labels. To mitigate this issue, we combine VoxCeleb2 with the LRS3 dataset, which contains manually annotated transcriptions corresponding to 433 hours of videos.
Since the LRS3 dataset does not include speaker ID information, we first generate the speaker ID information for both datasets using a method similar to the one described in (Kim, Kim, and Ro 2022), employing a face recognition model. Since there may be overlapping speakers between the two datasets, we take the following approach: For each video in LRS3 and VoxCeleb2, we divide the video into three equal segments and extract one face image from each segment. Then, we employ the Facenet (Serengil and Ozpinar 2024) face recognition model to encode the facial images into features of speaker information. To identify overlapping speakers between the two datasets, we compare the feature vectors from LRS3 and VoxCeleb2. This comparison is done using a similarity measure, such as cosine similarity, to identify the overlapping speakers.
@@
VoxLRS-SA is divided into two parts corresponding to each purpose. Firstly, the Baseline part of VoxLRS-SA is designed to train the baseline lip reading model. This training dataset is composed of 1.7K hours of video data from about 9K speakers. Secondly, the Adaptation part is proposed to adapt the pre-trained lip reading model to the target speaker and evaluate the effectiveness of the adaptation method. For this purpose, we construct the Adaptation part with 20 speakers, which do not overlap with the 9K speakers of the Baseline part. The train split contains video data varying from 45 minutes per speaker, and the test and val sets contain about 5 minutes of data for each of the 20 speakers. The detailed data information is shown in Table 2.
VoxCeleb2 (Chung, Nagrani, and Zisserman 2018) is a speaker verification dataset. It contains only speaker information for over 6,000 speakers and 2,442 hours of multilingual videos. By referring to the (Shi et al. 2022), we utilize an English portion, which amounts to 1,326 hours.
LRS3 (Afouras, Chung, and Zisserman 2018) is a widely used sentence-level lip reading dataset comprising 433 hours of English talking face videos and human-annotated transcriptions. These talking face videos are collected from TED and TEDx talks, thus contain diverse poses.